OpenStudy (anonymous):

How do I find an equation in standard form of the line parallel to the line with the equation -3x + 8y =10

OpenStudy (anonymous):

y = (3/8)x + C

OpenStudy (anonymous):

Ax + By = C is standard form

OpenStudy (anonymous):

hey toobulkeh

OpenStudy (anonymous):

since it is parallel then : -3/8 so the slope is -3/8

OpenStudy (anonymous):

hey kevin right rook, so -3 A, 8 is B and 10 is C. If you adjust C, the line moves up and down the graph, and retains the same behavior (slope) as your original equation.

OpenStudy (anonymous):

the slope of ur equation is 3/8 , so if u want to have a line parallel, as long as ur slope is 3/8, it does not matter what ur C is, in this case, the y-intercept 10 can be any value

OpenStudy (anonymous):

Rook, please reply here instead of direct messaging me.

OpenStudy (anonymous):

K / how do I do the sameequation -3x +8y =10 and passing through the point (4,0)

OpenStudy (anonymous):

a point is a coordinate pair, (x,y), so you can set X = 4 and Y = 0 and see what the value is for C that you get.

OpenStudy (anonymous):

-3 (4) + 8(0) = C

OpenStudy (anonymous):

Then just re-write it by using that C in your original equation.

OpenStudy (anonymous):

Thank you for the help !